# Atmospheric Density-Compensating Model Predictive Control for Targeted Reentry of Drag-Modulated Spacecraft

### v1

@@ 2.2 Drag Modulation
The atmospheric drag acting on a spacecraft in LEO causes its orbit to decay until it ultimately reenters the Earth’s atmosphere. The rate at which this decay occurs is related to the aerodynamic and mass properties of the spacecraft. A spacecraft that experiences a larger drag force relative to its mass will decay more quickly. The amount of drag compared to the spacecraft’s mass can be described by the ballistic coefficient, defined in this paper as
$C_{b}=\frac{C_{d}A}{2m},$ (8)
where $C_{d}$ is the drag coefficient of the spacecraft, $A$ is the cross-sectional area and $m$ is the mass of the spacecraft. The acceleration experienced by the spacecraft due to drag is written as
$a_{d}=\frac{\frac{1}{2}\rho v^{2}C_{d}A}{m},$ (9)
@@ 2.3 Atmospheric Density Prediction
Predicting the path that a spacecraft will follow under the influence of atmospheric drag requires advance knowledge of the atmospheric density along that path in order to compute the drag force. Models of the atmosphere are used to predict the density that a spacecraft will encounter. High-fidelity models such as NRLMSISE-00 [34] and Jacchia-type models [35] are able to account for variations in the atmospheric density due to latitude, longitude, altitude, time of day, day of year, as well as variations due to the space weather environment. While position and time are easily known along a given trajectory, the space weather environment is difficult to predict.
The state of Earth’s magnetic field as well as the activity of the Sun can cause the density of the atmosphere to vary by orders of magnitude, depending on the altitude. The effect of Earth’s magnetic field is included in high-fidelity models through the use of a geomagnetic index, such as $K_{p}$ or $a_{p}$, as an input. The activity of the Sun is included in atmosphere models using a solar index such as $F10.7a$ [36], which is measured by observing radio flux at a wavelength of $10.7$cm. While the effects of the geomagnetic index and solar flux are accounted for in high-fidelity atmosphere models, obtaining advance knowledge of atmospheric density from these models requires advance knowledge of the space weather environment, as described by these indices.
Drag-modulated spacecraft trajectories may have a duration ranging from several weeks to many months, requiring predictions of density along the entire trajectory. However, predicting the space weather environment is challenging and forecasts lose their accuracy after projecting only a few days into the future [37, 38], leading to poor accuracy in density predictions and errors in propagating the spacecraft state [36]. Additionally, even if perfect knowledge of the future space weather environment is available, the predictions made by these models still suffer from errors of $15-30\%$ RMS [39]. This uncertainty in the atmospheric density is a primary factor limiting the accuracy of orbital predictions [26].
@@ 3.1 Guidance Trajectory Generation
Guidance trajectories from the initial condition of the spacecraft to the desired latitude and longitude at the entry interface altitude are generated using a simplified method based on work by Omar [24, 27]. The majority of the guidance generation algorithm presented in this section is not considered a contribution of this work, but is performed in order to subject the proposed estimation and control methods to a variety of different trajectories and conditions. A small contribution of this paper is related to the choice of nominal ballistic coefficients when performing iteration the guidance trajectory in order to maximize the available range of ballistic coefficient for trajectory tracking. This minor contribution is discussed in Section 3.1.4. Those intending on implementing a drag-modulated reentry guidance algorithm are encouraged to review Refs. [24, 27].
The guidance scheme assumes that the spacecraft has an initial ballistic coefficient, $C_{b1}$, and that this ballistic coefficient is held constant until time $t_{\text{swap}}$ at which point the ballistic coefficient of the spacecraft changes to $C_{b2}$. The orbital decay of a spacecraft following this scheme is shown in Fig. 2. The guidance problem is then a matter of finding the values of the guidance parameters $C_{b1}$, $C_{b2}$ and $t_{\text{swap}}$ that result in proper entry interface targeting.
Figure 2: Orbital decay of spacecraft with a change in ballistic coefficient at a $t_{swap}$ of approximately 12 days.
The guidance algorithm makes use of analytical approximations that describe how a given, numerically propagated trajectory is changed by a change in the guidance parameters. These analytical equations are derived from
$\Delta\theta_{1}=\Delta\theta_{10}\frac{C_{b10}}{C_{b1}}$ (15)
and
$\Delta t_{1}=\Delta t_{10}\frac{C_{b10}}{C_{b1}}$ (16)
where $\Delta t_{10}$ and $\Delta\theta_{10}$ are the change in time and true anomaly during orbital decay between two semi-major axes for a spacecraft with a ballistic coefficient of $C_{b10}$. Assuming a circular orbit and an exponential atmospheric density model, these equations estimate that if the ballistic coefficient changes from $C_{b10}$ to $C_{b1}$ then the change in time and true anomaly will be $\Delta t_{1}$ and $\Delta\theta_{1}$ for the new trajectory. It is also shown that under these assumptions, that the average rate of change of true anomaly during the decay between two semi-major axes depends only of the initial and final semi-major axis and is independent of the ballistic coefficient. These analytical approximations are applied to different phases of the deorbit to derive relationships between the guidance parameters and the location of the spacecraft at the entry interface altitude.
The algorithm is initialized with an initial guess of the trajectory. Then the analytical approximations are used to compute new estimates of the guidance parameters that change the previously propagated trajectory in order to correct errors in the targeting. The trajectory is iteratively refined in distinct latitude and longitude targeting steps.
3.1.1 Algorithm Initialization
3.1.2 Latitude Targeting
An angle $\phi$ is defined in the orbital plane from the ascending node of the orbit to the position of the spacecraft. There are two angles, $\phi_{d1}$ and $\phi_{d2}$, that correspond to any desired final latitude. A numerically propagated trajectory will sweep out a total angle of $\Delta\phi_{t}$. The trajectory can be altered to land at the desired latitude by changing $\Delta\phi_{t}$ by a $\Delta\phi_{d}$ that is chosen to ensure that the spacecraft is at an angle of $\phi_{d1}$ or $\phi_{d2}$ at the final time. This results in a desired total swept angle of $\Delta\phi_{t,d}=\Delta\phi+\Delta\phi_{d}$. It is shown that this change in total angle can be obtained by a change in when the switch between ballistic coefficients occurs by
where $\omega_{2,avg}$ is an average angular velocity calculated from the numerically propagated orbit. A change in $\Delta\phi_{t}$ of an integer multiple of $2\pi$ radians results in the spacecraft reaching the same latitude, but a integer number or orbits earlier or later. As a result, if the spacecraft starts at a sufficiently high altitude, there are many possible values of $\Delta\phi_{d}$ that target the desired latitude leading to many feasible values of $t_{\text{swap}}$. If the initial altitude is too low, there may be no feasible choices of $t_{\text{swap}}$.
3.1.3 Longitude Targeting
After latitude targeting, the deorbit trajectory will sweep out an angle, $\Delta\phi_{t,d}$, that ensures the desired latitude is reached. However, because the latitude targeting step does not address longitude, there will be a residual longitude error that requires reduction. The longitude error, $e_{\text{long}}$ is estimated by using the value of $t_{\text{swap}}$ from latitude targeting along with (15) and (16) to estimate the new orbital lifetime. The new longitude after latitude targeting can then be predicted based on the final longitude of the numerically propagated trajectory, the change in orbital lifetime due to $\Delta t_{\text{swap}}$, the rotation rate of the Earth and the precession rate of the orbit. The longitude error is then computed based on the predicted final longitude of the altered trajectory. This longitude error is addressed by changing the orbital lifetime while preserving the total angle swept out during deorbit. This will ensure that the spacecraft lands at the same latitude, but at a different time. Because the spacecraft reaches entry interface at a different time, the Earth will be in a different orientation below the orbit of the spacecraft, which itself will have precessed by a different amount. Changing the orbital lifetime, $\Delta t_{t}$, by
$\Delta t_{d}=\frac{e_{\text{long}}}{\omega_{e}-\omega_{\text{RAAN}}},$ (18)
The change in orbital lifetime of $\Delta t_{d}$ is accomplished without changing $\Delta\phi_{t}$ through a change in $C_{b1}$ and $C_{b2}$ [27]. Given a numerically propagated trajectory with an initial ballistic coefficient of $C_{b10}$ that changes to $C_{b20}$ at $t_{\text{swap}}$, new ballistic coefficients are estimated that give a trajectory with $\Delta\phi_{t,d}$ and $\Delta t_{t,d}$ through
and
where $\Delta\theta_{10}$ and $\Delta\theta_{20}$ are the total amount of true anomaly swept through from $t=0$ to $t=t_{\text{swap}}$ and $t_{\text{swap}}$ to the final time of the numerically propagated trajectory, respectively. Similarly, $\Delta t_{10}$ and $\Delta t_{20}$ are the duration of the trajectory phases before and after $t_{\text{swap}}$ in the propagated trajectory.
Figure 3: Orbital decay of spacecraft with a change in ballistic coefficient at $t_{swap}$ at approximately 12 days.
These equations assume that the change between ballistic coefficients occurs at the same semi-major axis. However, because $C_{b1}$ has changed, the spacecraft will reach this semi-major axis at a different time. Therefore $t_{\text{swap}}$ must be changed to accomodate this change in $C_{b1}$. The new swap time is estimated as
$t_{\text{swap,new}}=t_{\text{swap,old}}\frac{C_{b10}}{C_{b1}}.$ (21)
These new values of the guidance parameters can be used to numerically propagate the trajectory of the spacecraft from the initial condition to obtain a new guess of the guidance trajectory, to serve as the basis of the next iteration.
3.1.4 Iteration
After the trajectory is simulated using the initial guess of the guidance parameters, the resulting trajectory is used to perform latitude targeting. Assuming that the spacecraft starts at a sufficiently high altitude, during latitude targeting there are many choices of $t_{\text{swap}}$ that target latitude. One approach is to choose the $t_{\text{swap}}$ that results in the smallest residual longitude error [24], however after longitude targeting, this results in guidance ballistic coefficients close to the initial guess, which is chosen to be near the extremes in order to increase the effect of changing $t_{\text{swap}}$ during latitude targeting, resulting in limited control authority remaining for trajectory tracking.
@@ 3.1.4 Iteration
Another approach is to choose the $t_{\text{swap}}$ that provides the largest amount of controllability margin about the desired longitude [40]. However, this approach tends to produce guidance ballistic coefficients toward the bottom of the feasible range of ballistic coefficients, limiting the amount of control authority in one direction. To illustrate this, the approach from [40] is used to generate 250 guidance trajectories with a feasible ballistic coefficient range of $C_{b,\text{min}}=0.025$ m2/kg to $C_{b,\text{max}}=0.1$ m2/kg. As shown in Fig. 3, the resulting guidance ballistic coefficients are clustered around $C_{b}=2\frac{C_{b1}C_{b2}}{C_{b1}+C_{b2}}=0.04$ m2/kg. With this value of the guidance ballistic coefficients, the controller is able to increase the ballistic coefficient, and therefore the drag, by 150%. However, the drag can only be decreased by 37.5%, leading to far more control authority to increase drag compared to reducing drag. This degrades the ability of the controller to track trajectories when the density is larger than expected. Omar [27] addresses this by only using a portion of the feasible range of $C_{b}$ when generating the guidance in order to reserve some of the range of $C_{b}$ for tracking. However, more control authority can be preserved if the guidance trajectory is generated in such a way to produce guidance ballistic coefficients in the middle of the feasible range.
@@
The ballistic coefficient in the middle of the feasible range, $C_{b,\text{mid}}$ is chosen such that a spacecraft with this ballistic coefficient can reduce or increase the drag acting on it by the same factor, that is,
@@
For each feasible $t_{\text{swap}}$, the residual longitude error and the $\Delta t_{t,d}$ that corrects for it is predicted. The $C_{b1}$ and $C_{b2}$ that produce the desired orbital lifetime are then computed with (19) and (20). The $t_{\text{swap}}$ is then selected according to
@@
This $t_{\text{swap}}$ and the corresponding $C_{b1}$ and $C_{b2}$ are then numerically simulated, producing a trajectory that is used as the basis of the next iteration. This process repeats until a trajectory is produced with a final targeting error of less than some tolerance or until a maximum number of iterations is reached. The position and velocity of the spacecraft along this trajectory, $\mbf{r}^{gp}_{ECI}(t)$ and $\mbf{v}^{gp/ECI}_{ECI}(t)$, serve as the guidance trajectory for the estimation and control algorithms to track while $C_{b1}$, $C_{b2}$ and $t_{\text{swap}}$ form the nominal control strategy where the guidance ballistic coefficient is
@@ 3.2 Equations of Motion
A linear model is required for both the EKF as well as the proposed MPC approach. It is assumed that at all times, the spacecraft is in a nearly circular orbit and that the distance between the spacecraft and the guidance position is small compared to the radius of the Earth. These assumptions enable the use of the linear model of relative motion developed by Schweighart and Sedwick [41]. This model is similar to the Clohessy-Wiltshire equations [33] but differs in that it captures some effects of the $J2$ perturbation.
In the linear model, the motion in the cross-track direction is decoupled from the motion in the radial and in-track directions. Because the control relies on changes in drag, which acts primarily opposite to the in-track direction, and the lack of coupling to the cross-track direction in the model, cross-track errors are uncontrollable. However, if the initial cross-track error is small and the error in the radial and in-track directions are kept sufficiently small, the orbit of the spacecraft will precess at a similar rate to the guidance orbit which prevents excessive growth in the cross-track errors.
@@
Because this position and velocity is relative to the guidance trajectory, this state vector also represents the tracking errors. In the absence of density prediction error and control input, the rate of change of the state vector is found through
@@ 3.3 Density Prediction Error Estimation
A discrete-time extended Kalman filter is used to estimate the density prediction error, $\Delta\rho$, as well as the radial and along-track position and velocity of the spacecraft relative to its guidance trajectory as a byproduct. It is assumed that a GPS receiver is used to provide measurements of the position and velocity of the spacecraft relative to the guidance position and velocity. An estimate of the density error and the spacecraft state is made at the beginning of each time step, which is used within the MPC strategy.
@@ 3.3.1 Process Model
The state of the spacecraft is then augmented with the fraction $c_{\Delta\rho}$, which represents a normalized density prediction error, to form the state of the extended Kalman filter, given by
@@
The process that governs $\mbf{x}_{\text{dyn}}$ is given by (35). However, the process that governs $c_{\Delta\rho}$ can not be easily represented as a linear function of the spacecraft state, $\mbf{x}_{\text{dyn}}$. The evolution of $c_{\Delta\rho}$ is the result of complex interactions between the activity of the sun, the state of Earth’s magnetic field, the altitude of the spacecraft and many more factors that affect the density of Earth’s atmosphere. Due to the intractability of developing a linear process model, the evolution of $c_{\Delta\rho}$ is modeled as a random walk in discrete time.
@@
where $\mbf{L}=\begin{bmatrix}\mbf{0}_{3\times 3}&\mbf{1}_{3\times 3}\end{bmatrix}^{\mathsf{T}}$. The ZOH approximates all time-varying parameters in the model as constant thoughout each time step. The discrete-time model is recomputed at each time step using the values of the parameters at the beginning for the time step (e.g., $v(t)=v(t_{k})$ for $t\in\left(t_{k},t_{k+1}\right]$). The ZOH also accurately reflects that MPC holds the control input $\Delta C_{b}$ constant over each time step. The MATLAB function c2d.m is used to perform the discretization. The process noise, $\mbf{w}_{k}\in\mathbb{R}^{3}$, with a covariance of $\mbf{W}=\mathbb{E}\{\mbf{w}_{k}\mbf{w}_{k}^{\mathsf{T}}\}$, is added to the velocity states as well as $c_{\Delta\rho}$ to account for nonlinearities in the dynamics that are neglected by the model and to enable the estimate of $c_{\Delta\rho}$ to change over time.

### v2

@@ 2.2 Drag Modulation
The atmospheric drag acting on a spacecraft in LEO causes its orbit to decay until it ultimately reenters the Earth’s atmosphere. The rate at which this decay occurs is related to the aerodynamic and mass properties of the spacecraft. A spacecraft that experiences a larger drag force relative to its mass will have its orbit decay more quickly. The amount of drag compared to the spacecraft’s mass can be described by the ballistic coefficient, defined in this paper as
$C_{b}=\frac{C_{d}A_{\text{ref}}}{2m},$ (6)
where $C_{d}$ is the drag coefficient of the spacecraft, $A_{\text{ref}}$ is the cross-sectional area and $m$ is the mass of the spacecraft. This definition of the ballistic coefficient is inverted compared to the more typical definition and contains an additional factor of one half. Using this definition, the drag acceleration becomes proportional to $C_{b}$ which results in dynamic equations that are more convenient for the purposes of estimation and control. The acceleration experienced by the spacecraft due to drag is written as
$a_{d}=\frac{\frac{1}{2}\rho v^{2}C_{d}A_{\text{ref}}}{m},$ (7)
@@ 2.3 Atmospheric Density Prediction
Predicting the path that a spacecraft will follow under the influence of atmospheric drag requires advance knowledge of the atmospheric density along that path in order to compute the drag force. High-fidelity models of the atmosphere, such as NRLMSISE-00 [40] and Jacchia-type models [41], are used to predict the density that a spacecraft will encounter while accounting for variations in the atmospheric density due to latitude, longitude, altitude, time of day, day of year, as well as variations due to the space weather environment. While atmospheric density model inputs such as position and time are precisely known along a given trajectory, the space weather environment is difficult to predict accurately which results in a large source of uncertainty in the density.
The state of Earth’s magnetic field as well as the activity of the Sun can cause the density of the upper atmosphere to vary by orders of magnitude. The effect of Earth’s magnetic field is included in high-fidelity models through the $K_{p}$ and $a_{p}$ geomagnetic indices [42]. The activity of the Sun is included in atmosphere models using a solar index such as $F10.7a$ [42], which is measured by observing radio flux at a wavelength of $10.7$cm. While the effects of the geomagnetic index and solar flux are accounted for in high-fidelity atmosphere models, obtaining advance knowledge of atmospheric density from these models requires advance knowledge of the space weather environment, as described by these indices.
Drag-modulated spacecraft trajectories may have a duration ranging from several weeks to many months, requiring predictions of density along the entire trajectory. However, predicting the space weather environment is challenging and forecasts lose their accuracy after projecting only a few days into the future [43, 44], leading to poor accuracy in density predictions and errors in propagating the spacecraft state [42]. Additionally, even if perfect knowledge of the future space weather environment is available, the predictions made by these models still suffer from root mean square errors of up to $30\%$ over a trajectory [45, 46, 47]. This uncertainty in the atmospheric density is a primary factor limiting the accuracy of orbital predictions [28].
@@ 3.1 Guidance Trajectory Generation
Guidance trajectories are generated with a simplified method based on work by Omar [26, 29] from the initial condition of the spacecraft to the desired latitude and longitude at the entry interface altitude. The guidance scheme considers a single modulation of the ballistic coefficient from an initial value of $C_{b1}$ to a final value of $C_{b2}$ that occurs at a time $t_{\text{swap}}.$ The guidance trajectory generation problem then consists of finding the values of these guidance parameters that yield a trajectory with the desired latitude and longitude at the entry interface altitude.
Omar’s method involves deriving simplified analytical relationships between the guidance parameters and the entry interface location. Given an orbital decay trajectory, these equations predict how the time of flight and number of orbits made during the trajectory will change based on a change in the guidance parameters. These equations are used to compute new values of the guidance parameters that correct for the targeting errors of a given trajectory, enabling an initial guess of the guidance trajectory to be iteratively refined until proper targeting is achieved.
The iteration involves distinct latitude and longitude targeting steps. In the latitude targeting step, $t_{\text{swap}}$ is altered to change the total number of orbits made during the trajectory, affecting the final latitude. After $t_{\text{swap}}$ is obtained, $C_{b1}$ and $C_{b2}$ are computed to change the time of flight, which alters the final longitude as the orbit precesses and the Earth rotates. The orbital decay of a spacecraft following this scheme is shown in Fig. 2. A detailed summary of Omar’s algorithm adapted for use in this work is presented in the Appendix for completeness.
@@
If the spacecraft has sufficient control authority and the trajectory begins at a sufficiently high altitude, there may be many valid combinations of the guidance parameters that yield proper reentry targeting. One approach involves selecting the $t_{\text{swap}}$ that targets the desired latitude at entry interface with the smallest residual longitude error [26] to be corrected in the longitude targeting step. Another approach is to choose the $t_{\text{swap}}$ that provides the largest amount of controllability margin about the desired longitude [48]. However, this approach tends to produce guidance ballistic coefficients $C_{b1}$ and $C_{b2}$ toward the bottom of the feasible range of ballistic coefficients, limiting the amount of control authority in one direction. To illustrate this, the approach from [48] is used to generate 250 guidance trajectories with a feasible ballistic coefficient range of $C_{b,\text{min}}=0.025$ m2/kg to $C_{b,\text{max}}=0.1$ m2/kg. As shown in Fig. 3, the resulting guidance ballistic coefficients are clustered around $C_{b}=2\frac{C_{b,\text{min}}C_{b,\text{max}}}{C_{b,\text{min}}+C_{b,\text{max}}}=0.04$ m2/kg. With this value of the guidance ballistic coefficients, the controller is able to increase the ballistic coefficient, and therefore the drag, by 150%. However, the drag can only be decreased by 37.5%, leading to far more control authority to increase drag compared to reducing drag. This degrades the ability of the controller to track trajectories when the density is larger than expected. Omar [29] addresses this by only using a portion of the feasible range of $C_{b}$ when generating the guidance in order to reserve some of the range of $C_{b}$ for tracking. However, more control authority can be preserved if the guidance trajectory is generated in such a way to produce guidance ballistic coefficients in the middle of the feasible range.
@@
A minor contribution of this paper is to choose a ballistic coefficient in the middle of the feasible range, $C_{b,\text{mid}}$ such that a spacecraft with this ballistic coefficient can reduce or increase the drag acting on it by the same factor, that is,
@@
For each feasible $t_{\text{swap}}$, the $C_{b1}$ and $C_{b2}$ that produce the desired entry interface location are computed. The $t_{\text{swap}}$ is then selected according to
@@
The results presented in Section 4.4 demonstrate that this choice of $t_{\text{swap}}$ produces nominal ballistic coefficients that are closer to the center of the feasible range compared to Fig. 3.
Figure 2: Orbital decays of spacecraft with (a) constant ballistic coefficients and (b) a change in ballistic coefficient at a $t_{swap}$ of approximately 12 days.
Figure 3: Nominal ballistic coefficients for 250 trajectories with randomly sampled initial conditions and desired entry interface locations compared to the feasible range.
@@ 3.2 Equations of Motion
A linear model is required for both the EKF as well as the proposed MPC approach. It is assumed, at all times, the spacecraft is in a nearly circular orbit and that the distance between the spacecraft and the guidance position is small compared to the radius of the Earth. These assumptions enable the use of the linear model of relative motion developed by Schweighart and Sedwick [49]. This model is similar to the Clohessy-Wiltshire equations [39] but differs in that it captures some effects of the $J2$ perturbation.
In the linear model, the motion in the cross-track direction is decoupled from the motion in the radial and in-track directions. Cross-track errors are uncontrollable because the control input is based on changes in drag, which acts primarily opposite to the in-track direction, which is not coupled to the cross-track direction in the model. However, if the initial cross-track error is small and the error in the radial and in-track directions are kept sufficiently small, the orbit of the spacecraft will precess at a similar rate to the guidance orbit which prevents excessive growth in the cross-track errors.
@@
Because this position and velocity are relative to the guidance trajectory, this state vector also represents the tracking errors. In the absence of density prediction error and control input, the rate of change of the state vector is found through
@@ 3.3 Density Prediction Error Estimation
A discrete-time EKF is used to estimate the density prediction error, $\Delta\rho$, as well as the radial and along-track position and velocity of the spacecraft relative to its guidance trajectory as a byproduct. It is assumed that a GPS receiver is used to provide measurements of the position and velocity of the spacecraft relative to the guidance position and velocity. An estimate of the density error and the spacecraft state is made at the beginning of each time step, which is used within the MPC strategy.
@@ 3.3.1 Process Model
The state of the spacecraft is then augmented with the fraction $c_{\Delta\rho}$, which represents a normalized density prediction error, to form the state of the EKF, given by
@@
The subscript $KF$ is included in the definition of this state to indicate that it is the state associated with the EKF. The process that governs $\mbf{x}_{\text{dyn}}$ is given by (25). However, the process that governs $c_{\Delta\rho}$ can not be easily represented as a linear function of the spacecraft state, $\mbf{x}_{\text{dyn}}$. The evolution of $c_{\Delta\rho}$ is the result of complex interactions between the activity of the sun, the state of Earth’s magnetic field, the altitude of the spacecraft and many more factors that affect the density of Earth’s atmosphere. Due to the intractability of developing a linear process model, the evolution of $c_{\Delta\rho}$ is modeled as a random walk in discrete time. Random walk processes have been used in the literature to estimate atmospheric density uncertainty [51, 52, 53], which provides additional confidence in this modeling choice.
@@
where $\mbf{L}=\begin{bmatrix}\mbf{0}_{3\times 3}&\mbf{1}_{3\times 3}\end{bmatrix}^{\mathsf{T}}$. The ZOH approximates all time-varying parameters in the model as constant thoughout each time step. The discrete-time model is recomputed at each time step using the values of the parameters at the beginning for the time step (e.g., $v(t)=v(t_{k})$ for $t\in\left(t_{k},t_{k+1}\right]$). The ZOH also accurately reflects that MPC holds the control input $\Delta C_{b}$ constant over each time step. The process noise, $\mbf{w}_{k}\in\mathbb{R}^{3}$, with a covariance of $\mbf{W}=\mathbb{E}\{\mbf{w}_{k}\mbf{w}_{k}^{\mathsf{T}}\}$, is added to the velocity states as well as $c_{\Delta\rho}$ to account for nonlinearities in the dynamics that are neglected by the model and to enable the estimate of $c_{\Delta\rho}$ to change over time.